It takes 35 minutes for 10 people to paint 7 walls.
How many minutes does it take 6 people to paint 6 walls?

Answer by Theo(13342)

r*p*t = q is the general formula that's used for this type of problem.

r = rate of one person
p = number of people
t = time
q = quantity

in your problem, you have:

p = 10
t = 35 minutes
q = 7 walls

r*p*t = q becomes r*10*35 = 7

solve for r to get r = 7/(10*35) = 7/350.

this means that each person can paint 7/350 of a wall in 1 minute.

now you have 6 people painting 6 walls and you want to know how long that will take if each person works at the same rate as before.

r*p*t = q becomes 7/350 * 6 * t = 6

solve for t to get t = 6 / (7/350 * 6) which becomes 6 * 350 / 7 / 6 which becomes 350 / 7 which becomes 50.

it will take 6 people 50 minutes to paint 6 walls if each person can paint an average of 7/350 of a wall in 1 minute.

r*p*t = q becomes 7/350 * 6 * 50 = 6
simplify to get 6 = 6.
this confirms that 50 minutes is the correct solution.